# NestedFP: High-Performance, Memory-Efficient Dual-Precision Floating Point Support for LLMs

### v1

@@ Appendix B Extended Kernel Performance for NestedFP16
Appendix C Extended End-to-End Throughput Evaluation
@@
Figure 10: Extended end-to-end throughput evaluation across different input/output token configurations. Each subplot shows throughput comparison for FP16, NestedFP16, FP8 and NestedFP8 across four models (Llama 3.1 8B, Mistral Nemo, Phi-4, Mistral Small) under different request size settings: (a) 32 input, 512 output tokens (b) 1024 input, 512 output tokens (c) 32 input, 32 output tokens and (d) 1024 input, 32 output tokens.
We extend our end-to-end evaluation to further demonstrate the robustness of NestedFP. Figure 10 presents results for four additional input/output configurations: (32, 512), (1024, 512), (32, 32), and (1024, 32). Across all settings, NestedFP16 exhibits minimal overhead compared to the FP16 baseline, with an average degradation of 3.88% and a maximum of 8.22%. On a per-model basis, the average overhead is 5.15% for LLaMA 3.1 (8B), 4.19% for Mistral-Nemo (12B), 3.35% for Mistral-Small (24B), and 2.81% for Phi-4 (14B). In contrast, NestedFP8 achieves significant speedups over NestedFP16, with average throughput gains of 1.26$\times$ for LLaMA 3.1, 1.33$\times$ for Mistral-Nemo, 1.53$\times$ for Mistral-Small, and 1.36$\times$ for Phi-4 across all input/output shapes. Notably, the largest model, Mistral-Small, benefits the most from the NestedFP8 quantization. These results indicate that the performance gains of NestedFP are consistent and robust across a wide range of input/output configurations.
When compared to Torch FP8, NestedFP8 maintains competitive performance, achieving 96.84% of Torch FP8’s throughput for LLaMA 3.1, 97.76% for Mistral-Nemo, 98.79% for Mistral-Small, and 96.77% for Phi-4. Our limited search space shows potential gain that can be achieved with further FP8 kernel optimization. We note that we did adopt Stream-K scheduling for NestedFP8.
Our evaluation setup mirrors the configuration used in the main throughput experiments. All tests were conducted on a single NVIDIA H100 GPU using the vLLM 0.8.3 V1 engine with PyTorch 2.6.0+cu12.4 for both the baseline and our proposed NestedFP implementations. Chunked prefill, enabled by default in the vLLM V1 engine, was used throughout the evaluation. This feature segments the prefill phase into smaller chunks, which are processed concurrently with decoding tokens within the same batch. At each iteration, the GEMM operation shape $(M,N,K)$ is determined based on the number of tokens grouped in that step, with $M$ denoting the total token count. To systematically evaluate throughput under different batch sizes, we adjusted the max_num_batched_token and max_num_seqs parameters to constrain the batch size accordingly.
@@ Appendix D CUTLASS FP16 GEMM Kernel
We extend the open-source NVIDIA CUTLASS 3.6 library to support our custom format. We begin by describing the efficient design of GEMM kernels within CUTLASS for the NVIDIA H100 GPU and introduce key components that define our design space. In the following section, we benchmark our CUTLASS kernels against PyTorch to validate strong performance of CUTLASS baseline.
D.1 GEMM Kernel Design
Figure 11: (a) Data parallel scheduling. A single output tile is processed by a single thread block. (b) Stream-K scheduling. A single output tile is processed collaboratively across multiple thread blocks. (c) DRAM to shared memory copy handled by 4 warps in producer warp group and TMA engine.
Figure 12: (a) Non-cooperative kernel. A single warp group is responsible for processing entire input tile. (b) Cooperative kernel. Two warp groups are each responsible for processing half of the first input tile and corresponding output tile.
GEMM operates on a weight matrix of shape $N\times K$ and an activation matrix of shape $M\times K$. To parallelize this computation, CUTLASS partitions the input matrices into tiles of size $T_{n}\times T_{k}$ and $T_{m}\times T_{k}$, respectively. These tiles define the granularity at which work is distributed across thread blocks.
As illustrated in Figure 11, each input tile is processed along the $K$ dimension to produce a corresponding output tile, which is accumulated in registers. CUTLASS employs a persistent tile scheduler, assigning each thread block to a fixed streaming multiprocessor (SM) for the duration of execution. In data-parallel scheduling, each SM processes the full $K$ dimension for a given output tile independently. In contrast, stream-K scheduling enables multiple thread blocks to collaboratively compute a single output tile by dividing the $K$-dimension workload across thread blocks, as shown in Figure 11. The partial results are later merged to form the final output. In all cases, the computed output tiles are written back to global memory (DRAM).
To hide DRAM latency and enable compute–memory overlap, tiles are asynchronously loaded from global memory into shared memory. As shown in Figure 11, shared memory is partitioned to hold multiple tiles, allowing future iterations to be staged while the current tile is being processed. On NVIDIA H100 GPUs, this transfer is managed by the Tensor Memory Accelerator (TMA) engine. CUTLASS uses warp specialization to pipeline these operations: one warp group (four warps) handles TMA data transfers (producers), while one or two other warp groups (consumers) execute matrix-multiply-accumulate (MMA) operations, depending on the kernel configuration.
Once data is available in shared memory, the kernel proceeds with MMA using one of two configurations. In the RS (register-shared) kernel, weights are loaded into registers and processed with activations from shared memory—this approach enables bit-level manipulations required by our custom format. In contrast, SS (shared-shared) kernels source both operands directly from shared memory. On Hopper, SM90 MMA instructions require the second operand to come from shared memory. MMA operations are executed at the warp group granularity, requiring all warps in a group to be ready before issuing the instruction. Consequently, SIMT-level manipulations on weights must complete across all four warps before any warp group MMA operation begins.
CUTLASS supports both non-cooperative and cooperative kernel modes. Non-cooperative kernels use a single consumer warp group, while cooperative kernels employ two consumer groups that share activations but process distinct weight tiles. The behavioral differences between these modes are illustrated in Figure 12.
D.2 CUTLASS FP16 Baseline and cuBLAS W16A16 Kernel Comparison
Figure 13: Performance comparison between PyTorch torch.matmul and our CUTLASS baseline across 14 unique $(N,K)$ GEMM shapes. Each subplot represents a different shape configuration, with batch size $(M)$ on the $x$-axis and latency on the $y$-axis.

### v2

@@ Appendix B Extended Kernel Performance for NestedFP
Appendix C Extended End-to-End Throughput Evaluation for NestedFP
@@
Figure 9: Extended end-to-end throughput evaluation across different input/output token configurations. Each subplot shows throughput comparison for FP16 and NestedFP across four models (Llama 3.1 8B, Mistral Nemo, Phi-4, Mistral Small) under different request size settings: (a) 32 input, 512 output tokens (b) 1024 input, 512 output tokens (c) 32 input, 32 output tokens and (d) 1024 input, 32 output tokens.
We extend our end-to-end evaluation to further demonstrate the robustness of NestedFP. Figure 9 presents results for four additional input/output configurations: (32, 512), (1024, 512), (32, 32), and (1024, 32). Across all settings, NestedFP exhibits minimal overhead compared to the FP16 baseline, with an average degradation of 5.52% and a maximum of 7.69%. On a per-model basis, the average overhead is 5.97% for LLaMA 3.1 (8B), 6.67% for Mistral-Nemo (12B), 5.51% for Phi-4 (14B), and 5.05% for Mistral-Small (24B).
Our evaluation setup mirrors the configuration used in the main throughput experiments. Throughput was computed as the total number of generated tokens divided by the end-to-end latency for each test. All tests were conducted on a single NVIDIA H100 GPU using the vLLM 0.8.5 V1 engine with PyTorch 2.7.0+cu12.6 for both the baseline and our proposed NestedFP implementations. Chunked prefill, enabled by default in the vLLM V1 engine, was used throughout the evaluation. This feature segments the prefill phase into smaller chunks, which are processed concurrently with decoding tokens within the same batch. At each iteration, the GEMM operation shape $(M,N,K)$ is determined based on the number of tokens grouped in that step, with $M$ denoting the total token count. We set the maximum number of batched tokens to 8192 for the throughput test.
@@ Appendix D CUTLASS FP16 GEMM Kernel
D.1 CUTLASS FP16 Baseline and PyTorch Kernel Comparison
Figure 10: Performance comparison between PyTorch kernel and our CUTLASS baseline across 14 unique $(N,K)$ GEMM shapes. Each subplot represents a different shape configuration, with batch size $(M)$ on the $x$-axis and latency on the $y$-axis.
@@ D.1 CUTLASS FP16 Baseline and PyTorch Kernel Comparison
D.2 Kernel Search Space
The kernel design space consists of two variants: non-cooperative and cooperative kernels. For non-cooperative kernels, we perform a grid search over tile dimensions with $T_{m}\in{16,32,64,128,256}$, $T_{n}\in{64,128,256}$, and $T_{k}\in{64,128,256}$. Cooperative kernels adopt larger tiles with $T_{n}\in{128,256}$, and configurations that fail to compile are excluded. Non-cooperative kernels use a cluster shape of $(1,1,1)$, whereas cooperative kernels use $(2,1,1)$. We employ the persistent scheduler for non-cooperative kernels, and both the persistent and Stream-K [25] schedulers for cooperative kernels.